the ratio of model Q and model T cell phones at a store is 4:5 if the store expects to sell 80 phones of these two models combined, how many model Q phones will most likely be sold?
Answer by stanbon(75887)

the ratio of model Q and model T cell phones at a store is 4:5 if the store expects to sell 80 phones of these two models combined, how many model Q phones will most likely be sold?
---------
4:5 is the same as 4x:5x where 4x is the number of Q phones.
----------------
Equation:
4x + 5x = 80
90x = 80
x = 8/9
------
Ans: # of Q phones = 4x = 4(8/9) = 32/9 = 3 5/9
-----------------
Cheers,
Stan H.
---------------